# Variance Stabilizing Transformations for Electricity Price Forecasting in Periods of Increased Volatility

### v1

@@ Abstract
Accurate day-ahead electricity price forecasts are critical for power system operation and market participation, yet growing renewable penetration and recent crises have caused unprecedented volatility that challenges standard models. This paper revisits variance-stabilizing transformations (VSTs) as a preprocessing tool by introducing a novel parametrization of the asinh transformation, systematically analyzing parameter sensitivity and calibration window size, and explicitly testing performance under volatile market regimes. Using data from Germany, Spain, and France over 2015–2024 with two model classes (NARX and LEAR), we show that VSTs substantially reduce forecast errors, with gains of up to 14.6% for LEAR and 8.7% for NARX relative to untransformed benchmarks. The new parametrized asinh consistently outperforms its standard form, while rolling averaging across transformations delivers the most robust improvements, reducing errors by up to 17.7%. Results demonstrate that VSTs are especially valuable in volatile regimes, making them a powerful tool for enhancing electricity price forecasting in today’s power markets.
@@
Electricity price forecasting, power market, probabilistic inputs, load, renewable energy sources
@@ 1 Introduction and motivation
Electricity price forecasting (EPF) is essential to modern power system operation (kat:zie:18; mac:uni:wer:23; tan:etal:23). Accurate day-ahead price forecasts support critical tasks such as unit commitment, reserve scheduling, demand bidding, and risk management for utilities and market participants alike (zha:etal:22; mac:ser:uni:23). As electricity markets become liberalized and more interconnected, reliable forecasting becomes even more essential for ensuring informed system operation and cost-effective market participation.
@@
At the same time, forecasting electricity prices has become substantially more difficult (lag:mar:des:wer:21). The rapid expansion of renewable energy sources, together with regulatory changes and geopolitical shocks, has led to unprecedented price volatility (gia:par:pel:16; mar:bel:ren:19). Day-ahead prices now exhibit statistical features that are especially challenging for forecasting: sharp spikes, heavy-tailed distributions, strong seasonal cycles, and—increasingly since the mid-2010s—frequent zero or even negative values. Such characteristics violate the assumptions underlying many classical time-series approaches and degrade their performance in exactly the conditions when accurate forecasts are most valuable.
@@
A promising method for improving forecast robustness in electricity markets is the use of variance-stabilizing transformations (VSTs). These transformations are applied to price series prior to model estimation to reduce the influence of extreme values—especially crucial given the frequent spikes in power prices (cia:mun:zar:22). By stabilizing volatility, VSTs improve the statistical properties of the data and enhance the reliability of forecasting models (jan:tru:wer:wol:13). Historically, the logarithmic transformation was the technique of choice for its simplicity and effectiveness in reducing scale effects. However, the widespread emergence of zero and negative day-ahead prices rendered the log transform infeasible, motivating the search for alternatives. Since the comprehensive evaluation by uni:wer:zie:18, which confirmed that well-chosen VSTs can markedly enhance forecast accuracy they have become widely used in EPF literature (for example mar:bel:ren:19; kat:zie:21; lag:mar:des:wer:21; gun:ugu:oks:23).
@@
uni:wer:zie:18 systematically assessed multiple transformations across 12 electricity markets and two model classes, showing that well-chosen VSTs can improve forecast accuracy. However, that study adopted fixed parameter values for each transformation, relied on a single calibration window size, and was based on data from a relatively stable period (2010–2016), predating the recent surge in market volatility. Similar framework with fixed parameter values was adopted in many paper that compared different transformation function in EPF context (uni:wer:18; shi:etal:21; cia:mun:zar:22)
@@
By addressing methodological limitations of earlier studies and aligning the analysis with recent more volatile electricity markets, this paper provides new evidence on when and how VSTs enhance electricity price forecasting, and demonstrates their practical value as a tool for improving decision-making in liberalized power systems.
@@ 2 Data
The German EPEX-DE market is one of the most widely studied markets for electricity price forecasting (cia:mun:zar:22; kie:kus:16; mac:22; nit:wer:23), owing to its central position in Europe, high trading liquidity, and complex price dynamics, including frequent negative prices. The market is also characterized by a significant proportion of intermittent renewables; in 2024, the installed capacity of wind and solar power farms accounted for 58% of total installed capacity. In addition to hourly day-ahead prices, we include day-ahead forecasts of electricity demand and renewable (wind and solar) generation obtained from the ENTSO-E Transparency Platform.
The French EPEX (studied for example in ali:rit:24; uni:wer:zie:18; shi:etal:21) is characterized by a unique generation mix dominated by nuclear power, but in recent years has also seen a gradual rise in renewable energy integration, particularly wind. It plays a key role in cross-border electricity flows within Western Europe (sze:bak:23) and is increasingly exposed to price volatility driven by both internal and regional dynamics. As in the case of Germany, we use hourly day-ahead prices and incorporate ENTSO-E-provided forecasts of electricity load and renewable generation to capture the growing impact of renewables on price formation.
Though less represented in the literature (for example cia:mar:nas:23; gar:car:san:15; uni:24; yan:22), the Spanish OMIE market has become increasingly important due to its rapidly changing generation mix. Between 2015 and 2024, Spain’s solar power installed capacity share quadrupled — from around 5% to over 20%. Similarly, for Spain, we include hourly day-ahead prices, along with ENTSO-E forecasts of load and renewable generation. This allows for consistent modeling inputs across all three markets.
All time series have been processed to an hourly resolution. To ensure consistency, adjustments were made to account for missing data and changes to daylight saving time. Specifically, missing timestamps (e.g., during the spring shift) were imputed using the average of the preceding or following hour or day. Duplicate timestamps (e.g., during the fall shift) were replaced by their mean.
@@ 2.1 Calibration window size
Previous studies (mar:ser:wer:18; ser:uni:wer:19) have shown that the choice of calibration window size is critical to electricity price forecasting and that there is no universally optimal length. Although longer windows generally reduce estimator variance and are advantageous under stationarity, electricity markets are highly non-stationary due to structural shifts, such as increasing renewable energy source (RES) penetration, and exogenous shocks, such as the pandemic caused by the COVID-19 virus and the energy crisis triggered by the war in Ukraine. These changes may make long historical windows less relevant to current price dynamics. Conversely, shorter windows can adapt more quickly to recent patterns, though they may lack sufficient data for stable model estimation.
@@ 3.1.1 NARX model
The first base model used in this study is a feedforward neural network within the NARX (nonlinear autoregression with exogenous variables) framework. This model is implemented in a series-parallel architecture (xie:tan:lia:2009). The model setup, including the selection of inputs and the network architecture, was adopted directly from lip:uni:25, where it served as the baseline for point forecasting. The network consists of a single hidden layer with five neurons (tanh activation) and a linear output layer (see Figure 2). The inputs include lagged electricity prices, daily price extremes, the most recent price at midnight, day-ahead forecasts of load and renewable energy source (RES) generation, commodity prices (coal, gas, oil, and EUAs), and weekday dummies — consistent with the best practices established in earlier studies (bil:gia:del:rav:23; che:uni:wer:25; zie:wer:18). A separate model is trained for each hour using a rolling window approach. To reduce estimation variance, an ensemble of ten independently trained NARX models is used for each forecast, and the final prediction is taken as the average across ensemble members.
@@ 3.1.2 LEAR model
The second forecasting approach used in this study is a high-dimensional linear model, often referred to in the EPF literature as the LEAR (LASSO Estimated AutoRegressive ) model. Originally introduced by uni:now:wer:16 and zie:16:TPWRS, and subsequently adopted in several studies (e.g., jed:lag:mar:wer:22; lag:mar:des:wer:21; zie:wer:18), the LEAR framework captures complex temporal and cross-hour dependencies by including a large number of predictors. Estimation is handled via LASSO regression (tib:96) to control for overfitting.
$p_{d,h}$ $=\sum_{i=1}^{24}\left(\beta_{i}p_{d-1,i}+\beta_{24+i}p_{d-7,i}\right)+\beta_{49}p_{d-1}^{\min}+\beta_{50}p_{d-1}^{\max}$
@@
$+\beta_{150}\text{EUA}_{d-2}^{\text{close}}+\sum_{i=1}^{7}\beta_{150+i}D_{i}+\varepsilon_{d,h}$ (1)
The model includes lagged prices for all 24 hours on days d1d-1 and d7d-7, daily price extremes ($p^{\min}_{d-1}$ and $p^{\max}_{d-1}$), and day-ahead forecasts of load, and RES generation for all 24 hours on day $d$. It also incorporates the most recent closing prices (from day d2d-2) of key commodities – coal, natural gas, crude oil, and EU carbon emission allowances – as well as a full set of seven weekday dummies. In total, the model includes over 150 regressors, enabling it to capture rich dependencies across time and variables. As with the NARX model, forecasts are generated independently for each hour of the day using a rolling window estimation approach.
@@ 3.2 Variance Stabilizing Transformations
This study focuses on four transformations: the inverse hyperbolic sine (asinh) with a tunable slope parameter; the Box-Cox transformation; the mirror log (mlog) transformation; and the probability integral transform (PIT) with a t-distribution. These choices are based on two criteria: (i) strong empirical performance in the benchmark comparison conducted by uni:wer:zie:18; uni:wer:18; cia:mun:zar:22 and (ii) continued relevance and adoption in the EPF literature.
@@
The asinh transformation was first introduced to the EPF literature by sch:11 as a way to stabilize variance without losing interpretability for small values. Since then it has been employed in multiple studies in combination with variety of models (zie:wer:18; shi:etal:21; gro:kac:kru:23; jan:25). Its appeal lies in its ability to behave linearly around the origin while still compressing large positive or negative values. In this work we extend the classical form by introducing a tunable slope parameter $c>0$:
@@
where $\lambda$ is a shape parameter controlling the degree of variance stabilization. In the EPF literature, Box–Cox has been widely applied, often with $\lambda=0.5$ as a default choice (cia:mun:zar:22; zak:etal:24; lip:uni:25). Here we broaden the analysis by considering $\lambda\in\{0.0,0.1,\dots,1.0\}$ to examine how the optimal setting depends on market conditions and calibration window length.
@@
The mirror logarithm was introduced in the EPF context by uni:wer:zie:18 as a robust alternative to the standard logarithm. It is a straightforward generalization of the log-transform with a mirror image of the logarithm for negative values. Subsequent studies have employed mlog in both point and probabilistic forecasting models (kat:zie:18; mar:bel:ren:19; leh:sch:her:22). The transformation is defined as:
$Y_{d,h}=\text{sgn}(p_{d,h})\left[\log\left(|p_{d,h}|+\tfrac{1}{c}\right)+\log(c)\right],$ (6)
@@
The PIT framework is grounded in transforming data via its empirical cumulative distribution function, followed by mapping through the inverse of a reference distribution. In EPF, the normal distribution has been the standard choice, with successful applications in both linear and nonlinear forecasting models (shi:etal:21; cia:mar:nas:23; mac:ser:uni:23; uni:24). However, normal-based PIT may not sufficiently accommodate heavy-tailed behavior in volatile markets. To address this, we adopt the inverse $t$-distribution with $\nu$ degrees of freedom:

### v2

@@ Abstract
Accurate day-ahead electricity price forecasts are critical for power system operation and market participation, yet growing renewable penetration and recent crises have caused unprecedented volatility that challenges standard models. This paper revisits variance stabilizing transformations (VSTs) as a preprocessing tool by introducing a novel parametrization of the asinh transformation, systematically analyzing parameter sensitivity and calibration window size, and explicitly testing performance under volatile market regimes. Using data from Germany, Spain, and France over 2015–2024 with two model classes (NARX and LEAR), we show that VSTs substantially reduce forecast errors, with gains of up to 14.6% for LEAR and 8.7% for NARX relative to untransformed benchmarks. The new parametrized asinh consistently outperforms its standard form, while rolling averaging across transformations delivers the most robust improvements, reducing errors by up to 17.7%. Results demonstrate that VSTs are especially valuable in volatile regimes, making them a powerful tool for enhancing electricity price forecasting in today’s power markets.
@@
Electricity price forecasting, power market, Variance Stabilizing Transformation, adaptive model averaging, increased volatility
journal: Electric Power Systems Research
@@ 1 Introduction and motivation
Electricity price forecasting (EPF) is essential to modern power system operation (mac:uni:wer:23; tan:etal:23). Accurate day-ahead price forecasts support critical tasks such as unit commitment, reserve scheduling, demand bidding, and risk management for utilities and market participants alike (zha:etal:22). As electricity markets become liberalized and more interconnected, reliable forecasting becomes even more essential for ensuring informed system operation and cost-effective market participation (mac:lip:uni:25).
Beyond market participation and bidding strategies, electricity price forecasting also supports short-term operational decision making in power systems. As discussed by wer:14, short-term price forecasts provide information relevant for operational planning under high volatility, where price spikes and scarcity events reflect underlying system stress. On the demand side, price forecasts are used to guide the operational scheduling of flexible assets such as energy storage systems, where charging and discharging decisions depend directly on expected price dynamics (chi:zam:zar:pal:18). Moreover, forecasting errors can lead to economically suboptimal operational schedules, particularly during volatile periods, highlighting the importance of forecast robustness beyond purely statistical accuracy (zar:can:bha:10).
@@
At the same time, forecasting electricity prices has become substantially more difficult (lag:mar:des:wer:21). The rapid expansion of renewable energy sources, together with regulatory changes and geopolitical shocks, has led to unprecedented price volatility (gia:par:pel:16; mar:bel:ren:19). Day-ahead prices now exhibit statistical features that are especially challenging for forecasting: sharp spikes, heavy-tailed distributions, strong seasonal cycles, and –increasingly since the mid-2010s – frequent zero or even negative values. Such characteristics violate the assumptions underlying many classical time-series approaches and degrade their performance in exactly the conditions when accurate forecasts are most valuable.
@@
A promising method for improving forecast robustness in electricity markets is the use of variance-stabilizing transformations (VSTs). These transformations are applied to price series prior to model estimation to reduce the influence of extreme values – especially crucial given the frequent spikes in power prices (cia:mun:zar:22). By stabilizing volatility, VSTs improve the statistical properties of the data and enhance the reliability of forecasting models (jan:tru:wer:wol:13). Historically, the logarithmic transformation was the most common choice due to its simplicity. However, the widespread emergence of zero and negative day-ahead prices rendered the log transform infeasible, motivating the search for alternatives. Since the comprehensive evaluation by uni:wer:zie:18, which confirmed that well-chosen VSTs can markedly enhance forecast accuracy they have become widely used in EPF literature (for example mar:bel:ren:19; kat:zie:21).
@@
uni:wer:zie:18 systematically assessed multiple transformations across 12 electricity markets and two model classes, showing that well-chosen VSTs can improve forecast accuracy. However, that study adopted fixed parameter values for each transformation, relied on a single calibration window size, and was based on data from a relatively stable period (2010–2016), predating the recent surge in market volatility. Similar framework with fixed parameter values was adopted in many paper that compared different transformation function in EPF context (shi:etal:21; cia:mun:zar:22)
jed:lag:mar:wer:22 review the evolution of electricity price forecasting methods, from classical time-series and regression models to regularized statistical learning approaches and, more recently, deep-learning and hybrid architectures. While early EPF relied primarily on autoregressive formulations (wer:14), regularization techniques such as LASSO enabled robust high-dimensional models including the LEAR framework (zie:16:TPWRS) Although deep-learning models have shown promising results, large-scale empirical evidence indicates that well-designed statistical learning models remain highly competitive at the day-ahead horizon (wer:14; lag:mar:des:wer:21; jed:lag:mar:wer:22) Since variance-stabilizing transformations operate at the preprocessing stage, they are applicable across different forecasting paradigms; however, their impact is expected to diminish as model nonlinearity increases, with highly nonlinear architectures such as deep neural networks potentially mitigating the influence of extreme observations intrinsically. Accordingly, this study focuses on model classes for which preprocessing plays a key role in estimation stability and forecast robustness.
@@ 2 Data
The German EPEX-DE market is one of the most widely studied markets for electricity price forecasting (zie:wer:18; cia:mun:zar:22), owing to its central position in Europe, high trading liquidity, and complex price dynamics, including frequent negative prices. The market is also characterized by a significant proportion of intermittent renewables; in 2024, the installed capacity of wind and solar power farms accounted for 58% of total installed capacity. In addition to hourly day-ahead prices, we include day-ahead forecasts of electricity demand and renewable (wind and solar) generation
@@
The French EPEX is characterized by a unique generation mix dominated by nuclear power, but in recent years has also seen a gradual rise in renewable energy integration, particularly wind. It plays a key role in cross-border electricity flows within Western Europe (sze:bak:23) and is increasingly exposed to price volatility driven by both internal and regional dynamics. As in the case of Germany, we use hourly day-ahead prices and incorporate TSO forecasts of electricity load and renewable generation to capture the growing impact of renewables on price formation.
Spanish OMIE market has become increasingly important due to its rapidly changing generation mix. Between 2015 and 2024, Spain’s solar power installed capacity share quadrupled — from around 5% to over 20%. Similarly, for Spain, we include hourly day-ahead prices, along with TSO forecasts of load and renewable generation. This allows for consistent modeling inputs across all three markets.
All the time series were obtained from a publicly available source, namely, the ENTSO-E Transparency Platform. (transparency.entsoe.eu/) [^1] The day-ahead electricity prices were downloaded from from transparency.entsoe.eu/market/energyPrices for corresponding bidding zones (BZN) FR for France, ES for Spain and the DE-AT-LU bid zone until 30.9.2018 and DE-LU afterwards () for Germany. Both day-ahead total load forecasts (transparency.entsoe.eu/market/energyPrices) and day-ahead wind and solar generation forecasts (transparency.entsoe.eu/generation/forecast/windAndSolar/) were downloaded for contries (CTY) Germany, France and Spain. All data were downloaded in hourly or 15-minutes resolution and the latter one have been processed to an hourly resolution. To ensure consistency, adjustments were made to account for missing data and changes to daylight saving time. Specifically, missing timestamps (e.g., during the spring shift) were imputed using the average of the preceding or following hour or day. Duplicate timestamps (e.g., during the fall shift) were replaced by their mean [^2] The codes to preprocess the data can be provided from authors.
Additionally, we have enriched the dataset with commodity market indicators, recognizing the influence of global energy markets on electricity prices. These indicators include the closing prices of coal (API2), natural gas (TTF), crude oil (Brent), and carbon emission allowances (EUA), sourced from Investing.com [^3] EUA: www.investing.com/commodities/carbon-emissions
Brent: www.investing.com/commodities/brent-oil-historical-data
Coal: www.investing.com/commodities/newcastle-coal-futures-historical-data
Gas: www.investing.com/commodities/dutch-ttf-gas-c1-futures-historical-data
@@ 2.1 Calibration window size
Previous studies (mar:ser:wer:18) have shown that the choice of calibration window size is critical to electricity price forecasting and that there is no universally optimal length. Although longer windows generally reduce estimator variance and are advantageous under stationarity, electricity markets are highly non-stationary due to structural shifts, such as increasing renewable energy source (RES) penetration, and exogenous shocks, such as the pandemic caused by the COVID-19 virus and the energy crisis triggered by the war in Ukraine. These changes may make long historical windows less relevant to current price dynamics. Conversely, shorter windows can adapt more quickly to recent patterns, though they may lack sufficient data for stable model estimation.
@@ 3.1 Base Models
The forecasting models considered in this study are used as established benchmarks (mac:lip:uni:25) to evaluate the impact of variance-stabilizing transformations rather than to introduce new modeling architectures. The comparison between NARX and LEAR is intended to assess the robustness of the proposed transformation and selection schemes across models with different structural characteristics. A detailed analysis of model complexity or interpretability is outside the scope of this paper.
@@ 3.1.1 NARX model
The first base model used in this study is a feedforward neural network within the NARX (nonlinear autoregression with exogenous variables) framework. This model is implemented in a series-parallel architecture (xie:tan:lia:2009). The model setup, including the selection of inputs and the network architecture, was adopted directly from lip:uni:25, where it served as the baseline for point forecasting. The network consists of a single hidden layer with five neurons and a subsequent linear layer producing a single-valued output (see figure 2) The state of the i-th neuron in the hidden layer can be expressed as:
$h_{i}=f(\sum_{j=1}^{n}x_{i}w^{(h)}_{ji}+b^{(h)}_{i})$,
where $x_{j}$ is the j-th regressor and $n$ is the number of regressors. $f(\cdot)$ is a nonlinear activation function, in this study selected to be a hyperbolic tangent function. The final output layer is a linear combination of the hidden-layer neuron states:
$\hat{y}=\sum_{i=1}^{5}h_{i}w^{(o)}_{i}+b^{(o)}.$
Model parameters are estimated in Matlab using the trainlm function, which implements the Levenberg–Marquardt optimization algorithm (hag:men:1994). To mitigate overfitting, early stopping is applied: training is halted when the validation error, computed using a 9:1 training-to-validation split, does not improve for 10 consecutive epochs.
The inputs include lagged electricity prices, daily price extremes, the most recent price at midnight, day-ahead forecasts of load and renewable energy source (RES) generation, commodity prices (coal, gas, oil, and EUAs), and weekday dummies — consistent with the best practices established in earlier studies (bil:gia:del:rav:23; che:uni:wer:25). A separate model is trained for each hour using a rolling window approach. To reduce estimation variance, an ensemble of ten independently trained NARX models is used for each forecast, and the final prediction is taken as the average across ensemble members.
@@ 3.1.2 LEAR model
The second forecasting approach used in this study is a high-dimensional linear model, often referred to in the EPF literature as the LEAR (LASSO Estimated AutoRegressive ) model. Originally introduced by uni:now:wer:16 and zie:16:TPWRS the LEAR framework captures complex temporal and cross-hour dependencies by including a large number of predictors. Estimation is handled via LASSO regression (tib:96) to control for overfitting. Prior to estimation, all explanatory variables are standardized to have a mean of zero and a variance of one, while the dependent variable remains on its original scale. The LASSO penalty parameter is selected from a grid of 100 values via 7-fold cross-validation within each calibration window.
The model includes lagged prices for all 24 hours on days d1d-1 and d7d-7, daily price extremes ($p^{\min}_{d-1}$ and $p^{\max}_{d-1}$), and day-ahead forecasts of load, and RES generation for all 24 hours on day $d$. It also incorporates the most recent closing prices (from day d2d-2) of key commodities – coal, natural gas, crude oil, and EU carbon emission allowances – as well as a full set of seven weekday dummies and non-penalized intercept. In total, the model includes over 150 regressors, enabling it to capture rich dependencies across time and variables. As with the NARX model, forecasts are generated independently for each hour of the day using a rolling window estimation approach.
$p_{d,h}$ $=\beta_{0}+\sum_{i=1}^{24}\left(\beta_{i}p_{d-1,i}+\beta_{24+i}p_{d-7,i}\right)+\beta_{49}p_{d-1}^{\min}+\beta_{50}p_{d-1}^{\max}$
@@
$+\beta_{150}\text{EUA}_{d-2}^{\text{close}}+\sum_{i=1}^{7}\beta_{150+i}D_{i}+\varepsilon_{d,h}.$ (1)
@@ 3.2 Variance Stabilizing Transformations
This study focuses on four transformations: the inverse hyperbolic sine (asinh) with a tunable slope parameter; the Box-Cox transformation; the mirror log (mlog) transformation; and the probability integral transform (PIT) with a t-distribution. These choices are based on two criteria: (i) strong empirical performance in the benchmark comparison conducted by uni:wer:zie:18; cia:mun:zar:22 and (ii) continued relevance and adoption in the EPF literature.
@@
The asinh transformation was first introduced to the EPF literature by sch:11 as a way to stabilize variance without losing interpretability for small values. Since then it has been employed in multiple studies in combination with variety of models (shi:etal:21; gro:kac:kru:23; jan:25). Its appeal lies in its ability to behave linearly around the origin while still compressing large positive or negative values. In this work we extend the classical form by introducing a tunable slope parameter $c>0$:
@@
where $\lambda$ is a shape parameter controlling the degree of variance stabilization. In the EPF literature, Box–Cox has been widely applied, often with $\lambda=0.5$ as a default choice (cia:mun:zar:22; lip:uni:25). Here we broaden the analysis by considering $\lambda\in\{0.0,0.1,\dots,1.0\}$ to examine how the optimal setting depends on market conditions and calibration window length.
@@
The mirror logarithm was introduced in the EPF context by uni:wer:zie:18 as a robust alternative to the standard logarithm. It is a straightforward generalization of the log-transform with a mirror image of the logarithm for negative values. Subsequent studies have employed mlog in both point and probabilistic forecasting models (kat:zie:18; leh:sch:her:22). The transformation is defined as:
$Y_{d,h}=\text{sgn}(p_{d,h})\left[\log\left(|p_{d,h}|+\tfrac{1}{c}\right)+\log(c)\right].$ (6)
@@
The PIT framework is grounded in transforming data via its empirical cumulative distribution function, followed by mapping through the inverse of a reference distribution. In EPF, the normal distribution has been the standard choice, with successful applications in both linear and nonlinear forecasting models (shi:etal:21; cia:mar:nas:23). However, normal-based PIT may not sufficiently accommodate heavy-tailed behavior in volatile markets. To address this, we adopt the inverse $t$-distribution with $\nu$ degrees of freedom: